How do you graph the equation: -5x+3y=2
How do you graph the equation: y=-7
How do you graph the equation: -3y-2x=9
How do you graph the equation: x=8
How do you graph the equation: 4y-6x=0
How do you graph the equation: 3=-x
How do you graph the equation: 2y+5=-3

Answer by Alan3354(69443)

How do you graph the equation: -5x+3y=2
It's a straight line so 2 points are needed.
Pick 2 values for x and find y.
Plot the 2 points.
Draw a line thru them.
---------------------------
How do you graph the equation: y=-7
Pick 2 values for x.
Plot (x,-7) and plot the 2 points.
Draw a line thru them.
--------------------------------
How do you graph the equation: -3y-2x=9
It's a straight line so 2 points are needed.
Pick 2 values for x and find y.
Plot the 2 points.
Draw a line thru them.
---------------------------
How do you graph the equation: x=8
Pick 2 values for y.
Plot (-3,y) and plot the 2 points.
Draw a line thru them.
---------------------------
How do you graph the equation: 4y-6x=0
It's a straight line so 2 points are needed.
Pick 2 values for x and find y.
Plot the 2 points.
Draw a line thru them.
---------------------------
How do you graph the equation: 3=-x
Pick 2 values for y.
Plot (-3,y) and plot the 2 points.
Draw a line thru them.
-------------
How do you graph the equation: 2y+5=-3
It's a straight line so 2 points are needed.
Pick 2 values for x and find y.
Plot the 2 points.
Draw a line thru them.
---------------------------